# Dimension conservation of harmonic measures in products of hyperbolic spaces

### v1

@@ Abstract.
We show that the harmonic measure on a product of boundaries satisfies dimension conservation for a random walk with non-elementary marginals on a countable group acting on a product of hyperbolic spaces under finite first moment condition.
@@ 1. Introduction
Let $\pi$ be a probability measure on ${\bm{\Gamma}}$ such that marginals $\mu$ and ${\mu^{\star}}$ are non-elementary, i.e., their supports generate non-elementary subgroups in $\Gamma$ and in ${\Gamma^{\star}}$ as groups respectively. For such a $\pi$, a harmonic measure $\nu_{\pi}$ is defined on the product of boundaries $\partial\Gamma\times\partial{\Gamma^{\star}}$ (cf. Section 2.1). It is a unique probability measure satisfying that
$\nu_{\pi}=\pi\ast\nu_{\pi},\quad\text{where $\pi\ast\nu_{\pi}=\sum_{\bm{x}\in\Gamma}\pi(\bm{x})\bm{x}\nu_{\pi}$ and $\bm{x}\nu_{\pi}:=\nu_{\pi}\circ\bm{x}^{-1}$}.$
@@
In the above, we consider the natural action of ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ on $\partial\Gamma\times\partial{\Gamma^{\star}}$. The harmonic measure $\nu_{\pi}$ has marginals $\nu_{\mu}$ and $\nu_{{\mu^{\star}}}$ on $\partial\Gamma$ and on $\partial{\Gamma^{\star}}$ respectively, and these are determined by $\mu$ and by ${\mu^{\star}}$. The measure $\nu_{\mu}$ (or $\nu_{{\mu^{\star}}}$) for a single hyperbolic group and its generalization has attracted intensive studies, including the dimension, e.g., [Kai98, Led01], more recently, [BHM11, HS17, Tan19, DY23]. However, the harmonic measure $\nu_{\pi}$ for a product group has been studied only in a few cases (see [Vol21] for a special case of products of hyperbolic free product groups). Dimensional properties of such a measure exhibit new features since it contains different behaviors depending on the factors. This manifests an additional difficulty, which arises in a higher rank setting—–in that case, the boundaries are assembled in further intricate ways such as flag varieties. See [KLP11] for a thorough discussion on this matter of subject, and recent works [Les21, Rap21, LL23]. For related results on ergodic invariant measures for affine iterated function systems, see e.g., [KP96, FH09, Fen23]. We study the harmonic measure $\nu_{\pi}$ itself and the conditional measure $\nu_{\pi}^{\eta}$ of $\nu_{\pi}$ on $\partial\Gamma\times\partial{\Gamma^{\star}}$ for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{\Gamma^{\star}}$. It is shown that these are exact dimensional under a finite first moment condition. The dimension formula is a sum of ratios asymptotic entropy over drift corrected with differential entropy. For the conditional measures, we provide a sufficient condition for the strict positivity of dimension.
@@
Let us consider left invariant hyperbolic metrics $d$ and $d^{\star}$ quasi-isometric to word metrics in $\Gamma$ and in ${\Gamma^{\star}}$, respectively. They induce quasi-metrics $q$ and ${q^{\star}}$ in the compactified spaces $\Gamma\cup\partial\Gamma$ and ${\Gamma^{\star}}\cup\partial{\Gamma^{\star}}$, respectively, and let $\bm{q}({\bm{\xi}}_{1},{\bm{\xi}}_{2}):=\max\{q(\xi_{1},\xi_{2}),{q^{\star}}(\eta_{1},\eta_{2})\}$ for ${\bm{\xi}}_{i}=(\xi_{i},\eta_{i})\in(\Gamma\cup\partial\Gamma)\times({\Gamma^{\star}}\cup\partial{\Gamma^{\star}})$ and $i=1,2$ (cf. Section 2.1). Let us assume that $\pi$ has a finite first moment, i.e., $\sum_{\bm{x}\in{\bm{\Gamma}}}\bm{d}(\operatorname{{\bf id}},\bm{x})<\infty$ for the identity element $\operatorname{{\bf id}}$ and $\bm{d}(\bm{x}_{1},\bm{x}_{2}):=\max\{d(x_{1},x_{2}),d^{\star}(y_{1},y_{2})\}$ for $\bm{x}_{i}=(x_{i},y_{i})\in{\bm{\Gamma}}$ and $i=1,2$. The asymptotic entropy $h(\pi)$ is defined as the limit for $n$-fold convolutions $\pi_{n}:=\pi^{\ast n}$,
@@
Let $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ be non-elementary hyperbolic groups, and $\pi$ be a probability measure on ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$ on $\Gamma$ and on Γsuperscriptnormal-${\Gamma^{\star}}$ respectively. Suppose that $l(\Gamma,\mu)\geq l({\Gamma^{\star}},{\mu^{\star}})$. Then the harmonic measure $\nu_{\pi}$ on $\left(\partial\Gamma\times\partial{\Gamma^{\star}},\bm{q}\right)$ is exact dimensional, i.e., for $\nu_{\pi}$-almost every ${\bm{\xi}}\in\partial\Gamma\times\partial{\Gamma^{\star}}$,
@@
Let $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ be non-elementary hyperbolic groups, and $\pi$ be a probability measure on ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$ on $\Gamma$ and on Γsuperscriptnormal-${\Gamma^{\star}}$ respectively. For $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{\Gamma^{\star}}$, the conditional measure $\nu_{\pi}^{\eta}$ is exact dimensional on $(\partial\Gamma\times\partial{\Gamma^{\star}},\bm{q})$, i.e., for $\nu_{\pi}^{\eta}$-almost every ${\bm{\xi}}\in\partial\Gamma\times\partial{\Gamma^{\star}}$,
@@
Theorem 1.2 is shown in Theorem 1.3 in a more general setting. Following Furstenberg [Fur08, Definition 3.1], we say that a Borel probability measure $\nu$ on a product of compact metric space ${\mathcal{M}}\times{\mathcal{M}}^{\star}$ satisfies dimension conservation if the following holds. Let us consider the pushforward $\nu^{\star}$ and a system of conditional measures $\nu^{\eta}$ for $\eta\in{\mathcal{M}}^{\star}$ of $\nu$ associated with the projection ${\mathcal{M}}\times{\mathcal{M}}^{\star}\to{\mathcal{M}}^{\star}$: the measures $\nu$ and $\nu^{\star}$ are exact dimensional with dimension $\dim\nu$ and $\dim\nu^{\star}$ respectively, for $\nu^{\star}$-almost every $\eta\in{\mathcal{M}}^{\star}$, conditional measures $\nu^{\eta}$ are exact dimensional with dimension $\dim\nu^{\eta}$, and
@@
In general, it holds that $h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})\leq h({\mu^{\star}})$, and the equality holds if and only if $(\partial{{\mathcal{X}}^{\star}},\nu_{\mu^{\star}})$ is a Poisson boundary for the pair $(\operatorname{Isom}{{\mathcal{X}}^{\star}},{\mu^{\star}})$ (cf. Section 2.3). Let $l({\mathcal{X}},\mu)$ be the drift associated with a ${\mu^{\star}}$-random walk on ${{\mathcal{X}}^{\star}}$. Theorem 1.2 is generalized in this setting.
@@
Let $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ be countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ respectively, and ${\bm{\Gamma}}:=\Gamma\times{\Gamma^{\star}}$. Further let us consider a probability measure $\pi$ on ${\bm{\Gamma}}$ of the following form: For some $\alpha\in(0,1]$,
@@
with non-elementary probability measures $\lambda$ and ${\lambda^{\star}}$ on $\Gamma$ and on Γsuperscriptnormal-${\Gamma^{\star}}$ respectively, and a probability measure $\pi_{0}$ on ${\bm{\Gamma}}$. It holds that $h(\pi)-h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})>0$, where ${\mu^{\star}}$ is the marginal of $\pi$ on Γsuperscriptnormal-${\Gamma^{\star}}$.
@@
where $\mu_{{\rm diag}}((x,x^{\star})):=\mu(x)$ if $x=x^{\star}$, and 00 if otherwise. The $\pi^{\rho}$-random walk on $\Gamma\times\Gamma$ appears in the study of noise sensitivity problem on groups [BB23, Tan22]. By Theorems 1.2 and 1.4 applied to the case when $\Gamma={\Gamma^{\star}}$ and $\mu={\mu^{\star}}$, it holds that for all $\rho\in(0,1]$,
@@
This reproduces [Tan22, Theorem 3.1].
@@
This example is not covered by Theorem 1.4. Suppose that ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ for two hyperbolic groups $\Gamma$ and ${\Gamma^{\star}}$, and that there exists a (proper) surjective homomorphism $\Pi:\Gamma\to{\Gamma^{\star}}$. Let $\Delta:\Gamma\to\Gamma\times{\Gamma^{\star}}$ be the diagonal embedding $\Delta(x)=(x,\Pi(x))$ for $x\in\Gamma$. For a non-elementary probability measure $\mu$ on $\Gamma$ with finite first moment, let $\pi:=\Delta_{\ast}\mu$ be the pushforward of $\pi$ by $\Delta$. In this case, marginals of $\pi$ are $\mu$ on $\Gamma$ and $\Pi_{\ast}\mu$ on ${\Gamma^{\star}}$ respectively, where $\Pi_{\ast}\mu$ is the pushforward of $\mu$ by $\Pi$. Applying to Theorem 1.2 with ${\mu^{\star}}=\Pi_{\ast}\mu$ shows that
@@
This follows since $h(\pi)=h(\mu)$ and $h(\partial{\Gamma^{\star}},\Pi_{\ast}\mu)=h(\Pi_{\ast}\mu)$. It holds that $h(\mu)=h(\Pi_{\ast}\mu)$ if $\Pi$ is an isomorphism, and it depends on $\Pi$ whether a strict inequality $h(\mu)>h(\Pi_{\ast}\mu)$ holds or not. As a simple explicit example, let $\Gamma=F_{m+1}$ and ${\Gamma^{\star}}=F_{m}$ be free groups of rank $m+1$ and $m$ respectively for $m\geq 2$, equipped with word metrics associated with free bases. Further let $\Delta:F_{m+1}\to F_{m}$ be a homomorphism defined by sending the free basis in $F_{m+1}$ to the free basis in $F_{m}$. For the uniform distribution $\mu$ on the symmetrized free basis in $F_{m+1}$, the induced distribution on $F_{m}$ defines a simple random walk on $F_{m}$ with holding probability $1/(m+1)$. A computation yields $l(F_{m+1},\mu)=m/(m+1)$, $l(F_{m},\Pi_{\ast}\mu)=(m-1)/(m+1)$,
@@ 2.2. Hausdorff dimensions
Let $({\mathcal{M}},q)$ be a compact metrizable space ${\mathcal{M}}$ with a quasi-metric $q$. It is basically intended $(\partial{\mathcal{X}},q)$ or $(\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}},\bm{q})$. For a set $E$ in ${\mathcal{M}}$, let $\dim E$ denote the Hausdorff dimension of $E$ with respect to the quasi-metric $q$. The definition is recalled briefly. Let $|E|:=\sup\{q(\xi,\eta)\ :\ \xi,\eta\in E\}$. For all $\alpha,\Delta\in{\mathbb{R}}_{+}$ with $\Delta>0$, let
@@
Let $({\mathcal{X}},d)$ and $(\partial{{\mathcal{X}}^{\star}},d^{\star})$ be hyperbolic metric spaces with bounded growth at some scale. There exists a constant $L\geq 1$ such that the following holds for every Borel probability measure $\nu$ on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ and for every Borel set $F$ in $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ with $\nu(F)>0$. For $\nu$-almost every ${\bm{\xi}}\in F$, there exists a constant $r({\bm{\xi}})>0$ such that for every $r\in(0,r({\bm{\xi}}))$,
@@
The assumption on $({\mathcal{X}},d)$ implies that for every $\alpha\in(0,1)$ there exists a bi-Lipschitz embedding $f$ from $(\partial{\mathcal{X}},q^{\alpha})$ to some finite dimensional standard Euclidean space $({\tt E}_{0},\|\cdot\|_{{\tt E}_{0}})$ (cf. [BS00, Theorem 9.2] and [Ass83, 2.6. Proposition]). More precisely, there exists a constant $L^{0}\geq 1$ such that for all $\xi,\eta\in\partial{\mathcal{X}}$,
@@
The map ${\bm{f}}$ is homeomorphic onto the image since $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ is compact. The product space ${\tt E}$ is endowed with the maximum norm $\|\cdot\|_{{\tt E}}$ of the factors. Let $B_{{\tt E}}(\bm{v},r)$ denote the ball in ${\tt E}$ with respect to the norm. It holds that

### v2

@@ Abstract.
We show that the harmonic measure on a product of boundaries satisfies dimension conservation for a random walk with non-elementary marginals on a countable group acting on a product of hyperbolic spaces under the finite first moment condition.
@@ 1. Introduction
Let $\pi$ be a probability measure on ${\bm{\Gamma}}$ such that marginals $\mu$ and ${\mu^{\star}}$ are non-elementary, i.e., their supports generate non-elementary subgroups in $\Gamma$ and in ${\Gamma^{\star}}$ as groups respectively. For such a $\pi$, a harmonic measure $\nu_{\pi}$ is defined on the product of boundaries $\partial\Gamma\times\partial{\Gamma^{\star}}$ (cf. Section 2.1). It is the unique probability measure satisfying that
$\nu_{\pi}=\pi\ast\nu_{\pi},\quad\text{where $\pi\ast\nu_{\pi}=\sum_{\bm{x}\in{\bm{\Gamma}}}\pi(\bm{x})\bm{x}\nu_{\pi}$ and $\bm{x}\nu_{\pi}:=\nu_{\pi}\circ\bm{x}^{-1}$}.$
@@
In the above, we consider the natural action of ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ on $\partial\Gamma\times\partial{\Gamma^{\star}}$. The harmonic measure $\nu_{\pi}$ has marginals $\nu_{\mu}$ and $\nu_{{\mu^{\star}}}$ on $\partial\Gamma$ and on $\partial{\Gamma^{\star}}$ respectively, and these are determined by $\mu$ and by ${\mu^{\star}}$. The measure $\nu_{\mu}$ (or $\nu_{{\mu^{\star}}}$) for a single hyperbolic group and its generalization has attracted intensive studies, including the dimension, e.g., [Kai98, Led01], more recently, [BHM11, HS17, Tan19, DY23]. However, the harmonic measure $\nu_{\pi}$ for a product group has been studied only in a few cases (see [Vol21] for a special case of products of hyperbolic free product groups). Dimensional properties of such a measure exhibit new features since it contains different behaviors depending on the factors. This manifests an additional difficulty, which arises in a higher rank setting—–in that case, the boundaries are assembled in further intricate ways such as flag varieties. See [KLP11] for a thorough discussion on this matter of subject, and recent works [Les21, Rap21, LL23]. For related results on ergodic invariant measures for affine iterated function systems, see e.g., [KP96, FH09, Fen23]. We study the harmonic measure $\nu_{\pi}$ itself and the conditional measure $\nu_{\pi}^{\eta}$ of $\nu_{\pi}$ on $\partial\Gamma\times\partial{\Gamma^{\star}}$ for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{\Gamma^{\star}}$. It is shown that these are exact dimensional under a finite first moment condition. The dimension formula is a sum of ratios of asymptotic entropy over drift corrected with differential entropy. For the conditional measures, we provide a sufficient condition for the strict positivity of dimension.
@@
Let us consider left invariant hyperbolic metrics $d$ and $d^{\star}$ quasi-isometric to word metrics in $\Gamma$ and in ${\Gamma^{\star}}$, respectively. They induce quasi-metrics $q$ and ${q^{\star}}$ in the compactified spaces $\Gamma\cup\partial\Gamma$ and ${\Gamma^{\star}}\cup\partial{\Gamma^{\star}}$, respectively, and let $\bm{q}({\bm{\xi}}_{1},{\bm{\xi}}_{2}):=\max\{q(\xi_{1},\xi_{2}),{q^{\star}}(\eta_{1},\eta_{2})\}$ for ${\bm{\xi}}_{i}=(\xi_{i},\eta_{i})\in(\Gamma\cup\partial\Gamma)\times({\Gamma^{\star}}\cup\partial{\Gamma^{\star}})$ and $i=1,2$ (cf. Section 2.1). Let us assume that $\pi$ has a finite first moment, i.e., $\sum_{\bm{x}\in{\bm{\Gamma}}}\bm{d}(\operatorname{{\bf id}},\bm{x})\pi(\bm{x})<\infty$ for the identity element $\operatorname{{\bf id}}$ and $\bm{d}(\bm{x}_{1},\bm{x}_{2}):=\max\{d(x_{1},x_{2}),d^{\star}(y_{1},y_{2})\}$ for $\bm{x}_{i}=(x_{i},y_{i})\in{\bm{\Gamma}}$ and $i=1,2$. The asymptotic entropy $h(\pi)$ is defined as the limit for $n$-fold convolutions $\pi_{n}:=\pi^{\ast n}$,
@@
Let $\Gamma$ and ${\Gamma^{\star}}$ be non-elementary hyperbolic groups, and $\pi$ be a probability measure on ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$ on $\Gamma$ and on ${\Gamma^{\star}}$ respectively. Suppose that $l(\Gamma,\mu)\geq l({\Gamma^{\star}},{\mu^{\star}})$. Then the harmonic measure $\nu_{\pi}$ on $\left(\partial\Gamma\times\partial{\Gamma^{\star}},\bm{q}\right)$ is exact dimensional, i.e., for $\nu_{\pi}$-almost every ${\bm{\xi}}\in\partial\Gamma\times\partial{\Gamma^{\star}}$,
@@
Let $\Gamma$ and ${\Gamma^{\star}}$ be non-elementary hyperbolic groups, and $\pi$ be a probability measure on ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$ on $\Gamma$ and on ${\Gamma^{\star}}$ respectively. For $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{\Gamma^{\star}}$, the conditional measure $\nu_{\pi}^{\eta}$ is exact dimensional on $(\partial\Gamma\times\partial{\Gamma^{\star}},\bm{q})$, i.e., for $\nu_{\pi}^{\eta}$-almost every ${\bm{\xi}}\in\partial\Gamma\times\partial{\Gamma^{\star}}$,
@@
Theorem 1.2 is shown in Theorem 1.3 in a more general setting. Following Furstenberg [Fur08, Definition 3.1], we say that a Borel probability measure $\nu$ on a product of compact metric spaces ${\mathcal{M}}\times{\mathcal{M}}^{\star}$ satisfies dimension conservation if the following holds. Let us consider the pushforward $\nu^{\star}$ and a system of conditional measures $\nu^{\eta}$ for $\eta\in{\mathcal{M}}^{\star}$ of $\nu$ associated with the projection ${\mathcal{M}}\times{\mathcal{M}}^{\star}\to{\mathcal{M}}^{\star}$: the measures $\nu$ and $\nu^{\star}$ are exact dimensional with dimension $\dim\nu$ and $\dim\nu^{\star}$ respectively, for $\nu^{\star}$-almost every $\eta\in{\mathcal{M}}^{\star}$, conditional measures $\nu^{\eta}$ are exact dimensional with dimension $\dim\nu^{\eta}$, and
@@
In general, it holds that $h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})\leq h({\mu^{\star}})$, and the equality holds if and only if $(\partial{{\mathcal{X}}^{\star}},\nu_{\mu^{\star}})$ is a Poisson boundary for the pair $(\operatorname{Isom}{{\mathcal{X}}^{\star}},{\mu^{\star}})$ (cf. Section 2.3). Let $l({\mathcal{X}},\mu)$ be the drift associated with a $\mu$-random walk on ${\mathcal{X}}$. Theorem 1.2 is generalized in this setting.
@@
Let $\Gamma$ and ${\Gamma^{\star}}$ be countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ respectively, and ${\bm{\Gamma}}:=\Gamma\times{\Gamma^{\star}}$. Further let us consider a probability measure $\pi$ on ${\bm{\Gamma}}$ of the following form: For some $\alpha\in(0,1]$,
@@
with non-elementary probability measures $\lambda$ and ${\lambda^{\star}}$ on $\Gamma$ and on ${\Gamma^{\star}}$ respectively, and a probability measure $\pi_{0}$ on ${\bm{\Gamma}}$. It holds that $h(\pi)-h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})>0$, where ${\mu^{\star}}$ is the marginal of $\pi$ on ${\Gamma^{\star}}$.
@@
where $\mu_{{\rm diag}}((x,x^{\star})):=\mu(x)$ if $x=x^{\star}$, and 00 if otherwise. The $\pi^{\rho}$-random walk on $\Gamma\times\Gamma$ appears in the study of noise sensitivity problem on groups [BB23, Tan24]. By Theorems 1.2 and 1.4 applied to the case when $\Gamma={\Gamma^{\star}}$ and $\mu={\mu^{\star}}$, it holds that for all $\rho\in(0,1]$,
@@
This reproduces [Tan24, Theorem 3.1].
@@
This example is not covered by Theorem 1.4. Suppose that ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ for two hyperbolic groups $\Gamma$ and ${\Gamma^{\star}}$, and that there exists a (non-injective) surjective homomorphism $\Pi:\Gamma\to{\Gamma^{\star}}$. Let $\Delta:\Gamma\to\Gamma\times{\Gamma^{\star}}$ be the diagonal embedding $\Delta(x)=(x,\Pi(x))$ for $x\in\Gamma$. For a non-elementary probability measure $\mu$ on $\Gamma$ with finite first moment, let $\pi:=\Delta_{\ast}\mu$ be the pushforward of $\mu$ by $\Delta$. In this case, marginals of $\pi$ are $\mu$ on $\Gamma$ and $\Pi_{\ast}\mu$ on ${\Gamma^{\star}}$ respectively, where $\Pi_{\ast}\mu$ is the pushforward of $\mu$ by $\Pi$. Applying to Theorem 1.2 with ${\mu^{\star}}=\Pi_{\ast}\mu$ shows that
@@
This follows since $h(\pi)=h(\mu)$ and $h(\partial{\Gamma^{\star}},\Pi_{\ast}\mu)=h(\Pi_{\ast}\mu)$. It holds that $h(\mu)=h(\Pi_{\ast}\mu)$ if $\Pi$ is an isomorphism, and it depends on $\Pi$ whether a strict inequality $h(\mu)>h(\Pi_{\ast}\mu)$ holds or not. As a simple explicit example, let $\Gamma=F_{m+1}$ and ${\Gamma^{\star}}=F_{m}$ be free groups of rank $m+1$ and $m$ respectively for $m\geq 2$, equipped with word metrics associated with free bases. Further let $\Pi:F_{m+1}\to F_{m}$ be a homomorphism sending $x_{i}$ to $y_{i}$, $i=1,\dots,m$, and sending $x_{m+1}$ to the identity, where $\{x_{1},\dots,x_{m+1}\}$ denotes the free basis in $F_{m+1}$ and $\{y_{1},\dots,y_{m}\}$ denotes the free basis in $F_{m}$. For the uniform distribution $\mu$ on the symmetrized free basis in $F_{m+1}$, the induced distribution on $F_{m}$ defines a simple random walk on $F_{m}$ with holding probability $1/(m+1)$. A computation yields $l(F_{m+1},\mu)=m/(m+1)$, $l(F_{m},\Pi_{\ast}\mu)=(m-1)/(m+1)$,
@@ 2.2. Hausdorff dimensions
Let $({\mathcal{M}},q)$ be a compact metrizable space ${\mathcal{M}}$ with a quasi-metric $q$. It is basically intended as $(\partial{\mathcal{X}},q)$ or $(\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}},\bm{q})$. For a set $E$ in ${\mathcal{M}}$, let $\dim E$ denote the Hausdorff dimension of $E$ with respect to the quasi-metric $q$. The definition is recalled briefly. Let $|E|:=\sup\{q(\xi,\eta)\ :\ \xi,\eta\in E\}$. For all $\alpha,\Delta\in{\mathbb{R}}_{+}$ with $\Delta>0$, let
@@
Let $({\mathcal{X}},d)$ and $({{\mathcal{X}}^{\star}},d^{\star})$ be hyperbolic metric spaces with bounded growth at some scale. There exists a constant $L\geq 1$ such that the following holds for every Borel probability measure $\nu$ on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ and for every Borel set $F$ in $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ with $\nu(F)>0$. For $\nu$-almost every ${\bm{\xi}}\in F$, there exists a constant $r({\bm{\xi}})>0$ such that for every $r\in(0,r({\bm{\xi}}))$,
@@
The assumption on $({\mathcal{X}},d)$ implies that for every $\alpha\in(0,1)$ there exists a bi-Lipschitz embedding $f$ from $(\partial{\mathcal{X}},q^{\alpha})$ to some finite dimensional standard Euclidean space $({\tt E^{0}},\|\cdot\|_{{\tt E^{0}}})$ (cf. [BS00, Theorem 9.2] and [Ass83, 2.6. Proposition]). More precisely, there exists a constant $L^{0}\geq 1$ such that for all $\xi,\eta\in\partial{\mathcal{X}}$,
@@
The map ${\bm{f}}$ is a homeomorphism onto its image since $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ is compact. The product space ${\tt E}$ is endowed with the maximum norm $\|\cdot\|_{{\tt E}}$ of the factors. Let $B_{{\tt E}}(\bm{v},r)$ denote the ball in ${\tt E}$ with respect to the norm. It holds that